solve the equations. I can't seem to get the right answer for a and C.
{-2+3C=a
{-a=2-4C

Answer by ikleyn(52799)

.

From the given equations, you have


    -2 + 3C = -(2 - 4C)


Simplify and find C


    -2 + 3C = -2 + 4C

         3C =      4C

         4C - 3C = 0

            C    = 0


Now, substitute C= 0 into the first equation and find  


    -2 + 3*0 = a

       -2    = a.


ANSWER.  a = -2;  C = 0.

Solved.